# TPT-Bench: A Large-Scale, Long-Term and Robot-Egocentric Dataset for Benchmarking Target Person Tracking

### v1

@@ Abstract
Tracking a target person from robot-egocentric views is crucial for developing autonomous robots that provide continuous personalized assistance or collaboration in Human-Robot Interaction (HRI) and Embodied AI. However, most existing target person tracking (TPT) benchmarks are limited to controlled laboratory environments with few distractions, clean backgrounds, and short-term occlusions. In this paper, we introduce a large-scale dataset designed for TPT in crowded and unstructured environments, demonstrated through a robot-person following task. The dataset is collected by a human pushing a sensor-equipped cart while following a target person, capturing human-like following behavior and emphasizing long-term tracking challenges, including frequent occlusions and the need for re-identification from numerous pedestrians. It includes multi-modal data streams, including odometry, 3D LiDAR, IMU, panoptic, and RGB-D images, along with exhaustively annotated 2D bounding boxes of the target person across 35 sequences, both indoors and outdoors. Using this dataset and visual annotations, we perform extensive experiments with existing TPT methods, offering a thorough analysis of their limitations and suggesting future research directions. Our dataset, code and video are available at https://medlartea.github.io/tpt-bench/
@@ I Introduction
To effectively perform Human-Robot Interaction (HRI) and Embodied AI tasks in environments with people, robots must recognize their human counterparts and adapt their actions to provide personalized assistance. The ability to continuously recognize and track a target person, known as target person tracking (TPT), is an egocentric perceptual task related to navigation that has recently gained significant attention in the robotics and HRI communities[1, 2, 3, 4, 5, 6, 7]. While these frameworks demonstrate efficient identification, they are often evaluated in controlled laboratory settings with minimal distractions, clean backgrounds, and short tracking durations.
@@
Datasets for evaluating TPT are limited. Chen et al.[8] introduced a public TPT dataset featuring challenging scenarios such as short-term occlusions by rapid multi-person crossings, illumination changes, and appearance variations. However, it lacks more challenging scenarios, such as occlusions and distractions caused by people with similar appearances. To address this gap, some methods have proposed custom-built datasets[4, 5, 6, 7], but their experimental settings are constrained, typically involving only a few distractors and being collected in controlled laboratory environments with clean backgrounds. Such small-scale and structured environments limit the development of TPT.
@@
TPT focuses on continuous tracking of a target person, with related datasets including multiple-object tracking (MOT)[9, 10, 11] and single-object tracking (SOT)[12, 13, 14]. While SOT datasets address challenges like occlusion and pose changes, they are designed for generic object tracking and lack a moving egocentric viewpoint for performing robot companionship in crowded and unstructured environments. Such a scenario involves uncontrolled natural conditions, especially frequent occlusion and obstruction. In addition, SOT datasets offer limited trajectories for people and short-duration recordings. MOT datasets focus on human trajectories in crowded scenes, but the recoding time for individual subjects is still limited.
@@
To address these gaps, we propose a TPT dataset that captures diverse scenarios from a moving egocentric perspective in crowded environments, highlighting the challenges of frequent occlusion and obstruction from numerous pedestrians. The dataset focuses on a realistic HRI task: robot person following[1] or assistive socially-aware navigation[2]. It is collected by a human pushing a sensor-equipped cart while following a target person, as shown in Fig. 2. This human-human interaction behavior offers valuable insights for developing human-like person-following methods[15]. The dataset contains 3.35 hours of multi-modal data, including robot odometry, 3D LiDAR, panoptic and RGBD images, IMU data, and annotated 2D bounding boxes of the target person across diverse crowded environments, such as schools, offices, markets, and metro stations, spanning 35 sequences.
@@
  • A large-scale, long-term, multi-modal, and egocentric TPT dataset highlighting long-term tracking challenges of frequent target disappearance in diverse, crowded environments.
  • A unified evaluation framework that measures a TPT algorithm’s ability to identify and track the true target person with accuracy and confidence.
  • A systematic analysis and comparison of existing TPT algorithms on the TPT-Bench, LaSOT[14], and Chen’s dataset[8], demonstrating the TPT-Bench is more challenging with the objective of long-term tracking.
II Related Work
@@
LaSOT[14] SOT 20 106.8 0.3 7.8 31 1.6
LTB50[16] SOT 20 106.3 9.1 0.8 97 4.9
MOT20[11] MOT 2257 20.5 1.5 126 1617 0.7
JRDB[17] MOT 1805 65.7 11.5 47 7265 4.0
Chen’s dataset[8] TPT 11 340.5 1.8 3 4 0.4
TPT-Bench (ours) TPT 35 394.3 83.5 10.2 1357 38.8
TABLE I: Public TPT datasets and related person-tracking datasets, including MOT and SOT datasets. For SOT datasets, we focus solely on video sequences involving the person category. In the case of MOT datasets, which feature trajectories of multiple individuals, we report attributes for each person as a sequence across all videos. Therefore, MOT20[11] and JRDB[17] report 2257 and 1805 instances, respectively, but actually encompass only 4 and 27 scenarios. Typically, JRDB has only 27 scenarios labels open sourced, and only 16 contain robot-moving behavior. Comparison of long-term properties: the number of sequences (#sequences), the number of scenarios (#scenarios), the average length of a person (APL, s), the average disappearance length (ADL, s), the average distractors per second in the disappearance interval (ADNDS), the number of target disappearances (DSP), the average number of disappearances in sequence (ADN)$=$DSP///#sequences.
II-A Target Person Tracking
@@
Target person tracking aims to identify and track a specific person, which is crucial for personalized HRI scenarios where the robot needs to engage with its user over extended periods[1, 2]. Earlier TPT approaches often relied on visual tracking techniques, such as MOT or SOT, to track the target person. SOT-assisted methods initially identify the target person by selecting their bounding box in the first frame, then use an appearance model to search for the target within a nearby region[8, 18]. SOT techniques[19] aim to track an arbitrary object by simultaneously classifying and localizing it within a search region. However, due to the lack of prior knowledge about human appearance (e.g., from pre-trained models of people detection and person re-identification), these methods can easily track a wrong person in crowded environments, particularly after partial or total occlusions.
@@
Conversely, MOT techniques[20] assume prior knowledge of the tracked object, typically people. MOT techniques usually detect individuals using well-trained object detectors and track multiple people within the sensor’s field of view (FoV). Based on these well-tracked people, MOT-assisted TPT methods then identify the target person by selecting their identity [21, 22]. Such a solution is robust to short-term occlusions but still struggles with long-term occlusions since they often lack a target re-identification (ReID) capability after occlusions.
@@
To address these gaps, recent TPT methods typically combine MOT with a target ReID module. The MOT front-end serves two main purposes: 1) tracking people’s trajectories is essential for socially-aware path planning[22, 23, 24], which helps the robot assist the target person while respecting proxemics with others, and 2) by re-identifying the target from tracked individuals, established person ReID techniques[25] can be utilized in the back-end module to recognize the target person after partial or total occlusions. With the advancement of deep learning, deep-learning-based (DL-based) ReID features have become dominant due to their robustness to viewpoint and illumination changes[25].
@@
To leverage the power of DL-based features for target ReID, several approaches have been proposed[5, 6, 7, 4]. For example, [7] introduced an online boosting classifier using low-level CNN features, while [5] represented the target feature as a distribution and updated it continuously with newly archived features using a damped exponential moving average. Moreover, [4] observed a domain shift in pre-trained ReID features—those from video-surveillance data often fail to generalize to practical person-following environments. To address this, they proposed using online continual learning techniques to optimize ReID features by collecting valuable observations.
@@
This paper evaluates the most recent target-ReID methods[5, 4] and popular SOT trackers on TPT-Bench, as well as other related datasets like LaSOT[14] and Chen’s dataset[8], which track a target person over extended periods. It highlights the more challenging long-term tracking scenarios presented by TPT-Bench.
II-B Datasets related to Target Person Tracking
II-B1 Datasets for Target Person Tracking
@@
Datasets for evaluating TPT are limited. The only public dataset[26] is collected by teleoperating the robot following the target person, involving challenging scenarios such as quick multi-people crossings, illumination changes, and appearance variations. However, this dataset lacks challenging scenarios requiring person ReID, such as occlusion and similar appearances of distracting people. To mitigate this limitation, some methods propose custom-built datasets[4, 5, 6, 7], but their experimental scenarios are limited, involving only a few distracting people.
@@
Moreover, ROREAS[27] and ROREAS+[28] propose TPT datasets involving scenarios of stroke patient assistance in their daily rehabilitation. Although these include challenges such as frequent occlusions by multiple people and lighting changes, their scenarios are limited to stroke patient assistance and are private due to data protection laws.
II-B2 Datasets for SOT and MOT
@@
Since TPT focuses on continuous online tracking of a target person, related datasets include MOT datasets[9, 10, 11] and SOT datasets[12, 13, 14]. SOT datasets, such as LaSOT[14] and TrackingNet[13], focus on tracking arbitrary objects, including people, across various scenarios like dance, surveillance, and sports. While these datasets present challenges like occlusion, cluttered backgrounds, and pose variations, they lack robot-egocentric sequences designed for personalized robot assistance. In such scenarios, the tracked person may frequently be occluded or disappear for long periods. Additionally, the robot’s perspective differs from that of surveillance or handheld cameras in that people are often partially observed with body parts, making it more challenging to track the person.
@@
MOT datasets, such as MOT20[11], are designed for multi-person tracking in crowded environments with frequent occlusions. However, they focus on surveillance footage, where occlusion patterns differ significantly from those seen in robot observations. These datasets also provide limited long-term tracking data for a target person. The JRDB dataset[17], the most relevant to our work, includes egocentric data with 3D LiDAR, IMU and visual data, collected by a teleoperated robot in crowded environments. However, it does not represent a long-term tracking scenario in the context of HRI. Moreover, it assumes people remain within the camera’s field of view (FoV) and does not address the need to re-identify individuals when they reappear after being occluded.
@@
A quantitative comparison of person-tracking datasets is presented in Table I, with most metrics derived from a long-term tracking benchmark[16]. The TPT-Bench captures long-term robot companionship from an egocentric perspective in crowded, unstructured environments, resulting in more occlusions and longer disappearance periods compared to existing person-tracking datasets. Additionally, when a target reappears, its appearance may differ from its previous state in terms of pose, lighting, and scale, as shown in Figure 1, which presents additional challenges for target ReID. These difficulties are demonstrated in the experiments through a comprehensive comparison of current TPT algorithms on the TPT-Bench.
III The TPT-Bench Dataset
@@
In this section, we first describe the data collection platform and introduce the scenarios collected in this dataset. We then provide data statistics for the whole dataset, highlighting unique challenges present by the proposed dataset.
III-A Collecting Data
III-A1 Platform

### v2

@@ Abstract
Tracking a target person from robot-egocentric views is crucial for developing autonomous robots that provide continuous personalized assistance or collaboration in Human-Robot Interaction (HRI) and Embodied AI. However, most existing target person tracking (TPT) benchmarks are limited to controlled laboratory environments with few distractions, clean backgrounds, and short-term occlusions. In this paper, we introduce a large-scale dataset designed for TPT in crowded and unstructured environments, demonstrated through a robot-person following task. The dataset is collected by a human pushing a sensor-equipped cart while following a target person, capturing human-like following behavior and emphasizing long-term tracking challenges, including frequent occlusions and the need for re-identification from numerous pedestrians. It includes multi-modal data streams, including odometry, 3D LiDAR, IMU, panoramic images, and RGB-D images, along with exhaustively annotated 2D bounding boxes of the target person across 48 sequences, both indoors and outdoors. Using this dataset and visual annotations, we perform extensive experiments with existing SOTA TPT methods, offering a thorough analysis of their limitations and suggesting future research directions. Our dataset, code and video are available at https://medlartea.github.io/tpt-bench/.
@@ 1 Introduction
To effectively perform Human-Robot Interaction (HRI) and Embodied AI tasks in environments with people, robots must recognize their human counterparts and adapt their actions to provide personalized assistance. The ability to continuously recognize and track a target person, known as target person tracking (TPT), is an egocentric perceptual task related to navigation that has recently gained significant attention in the robotics and HRI communities (Islam et al., 2019; Singamaneni et al., 2024; Siva and Zhang, 2022; Ye et al., 2023, 2024; Rollo et al., 2024; Eirale et al., 2025; Koide et al., 2020; Zhao et al., 2024). While these frameworks demonstrate efficient identification, they are often evaluated in controlled laboratory settings with minimal distractions, clean backgrounds, and short tracking durations.
@@
Datasets for evaluating TPT are limited. Chen et al. (2017b) introduced a public TPT dataset featuring challenging scenarios such as short-term occlusions by rapid multi-person crossings, illumination changes, and appearance variations. However, it lacks more challenging scenarios, such as occlusions and distractions caused by people with similar appearances. To address this gap, some methods have proposed custom-built datasets (Ye et al., 2024; Rollo et al., 2024; Koide et al., 2020; Gross et al., 2017; Eisenbach et al., 2023), but their experimental settings are constrained, typically involving only a few distractors and being collected in controlled laboratory environments with clean backgrounds. Such small-scale and structured environments limit the development of TPT.
@@
TPT focuses on continuous tracking of a target person, with related datasets including multiple-object tracking (MOT) (Milan, 2016; Manen et al., 2017; Dendorfer, 2020) and single-object tracking (SOT) (Wu et al., 2013; Muller et al., 2018; Fan et al., 2021). While SOT datasets address challenges like occlusion and pose changes, they are designed for generic object tracking and lack a moving egocentric viewpoint for performing robot companionship in crowded and unstructured environments. Such a scenario involves uncontrolled natural conditions, especially frequent occlusion and obstruction. In addition, SOT datasets offer limited trajectories for people and short-duration recordings. MOT datasets focus on human trajectories in crowded scenes, but the recoding time for individual subjects is still limited.
@@
To address these gaps, we propose a TPT dataset that captures diverse scenarios from a moving egocentric perspective in crowded environments, highlighting the challenges of frequent occlusion and obstruction from numerous pedestrians. The dataset focuses on a realistic HRI task: robot person following (Islam et al., 2019) or assistive socially-aware navigation (Singamaneni et al., 2024). It is collected by a human pushing a sensor-equipped cart while following a target person, as shown in Figure 2. This human-human interaction behavior offers valuable insights for developing human-like person-following methods (Karunarathne et al., 2018). The dataset comprises 5.3 hours of multi-modal recordings, including robot odometry, 3D LiDAR, panoptic and RGB-D images, as well as IMU streams. In addition to the raw data, we provide frame-level 2D bounding box annotations of the target person, totaling 571,982 frames. Collected across 48 sequences in densely populated environments—such as schools, food stores, markets, plazas, and metro stations—the dataset enables long-term tracking episodes, with each sequence averaging 397.2 seconds in duration.
@@
  • A large-scale, long-term, multi-modal, and robot-egocentric TPT dataset highlighting long-term tracking challenges of frequent target disappearance in diverse, crowded environments.
  • An evaluation framework that measures a TPT algorithm’s ability to identify and track the true target person with accuracy and confidence.
  • A systematic analysis and comparison of existing TPT algorithms on the TPT-Bench, LaSOT (Fan et al., 2021), and Chen’s dataset (Chen et al., 2017b), demonstrating the TPT-Bench is more challenging with the objective of long-term tracking.
2 Related Work
@@
Table 1: Public TPT datasets and related person-tracking datasets, including MOT and SOT datasets. For SOT datasets, we focus solely on video sequences involving the person category. In the case of MOT datasets, which feature trajectories of multiple individuals, we report attributes for each person as a sequence across all videos. Therefore, MOT20 (Dendorfer, 2020) and JRDB (Martin-Martin et al., 2021) report 2257 and 1805 instances, respectively, but actually encompass only 4 and 27 scenarios. Typically, JRDB has only 27 scenarios labels open sourced, and only 16 contain robot-moving behavior. Comparison of long-term properties: the number of sequences (#sequences), the number of scenarios (#scenarios), the average length of a person (APL, s), the average disappearance length (ADL, s), the average distractors per second in the disappearance interval (ADNDS), the number of target disappearances (DSP), the average number of disappearances in sequence (ADN)$=$DSP///#sequences.
@@
LaSOT (Fan et al., 2021) SOT 20 106.8 0.3 7.8 31 1.6
LTB50 (Lukeźič et al., 2020) SOT 20 106.3 9.1 0.8 97 4.9
MOT20 (Dendorfer, 2020) MOT 2257 20.5 1.5 126 1617 0.7
JRDB (Martin-Martin et al., 2021) MOT 1805 65.7 11.5 47 7265 4.0
Chen’s dataset (Chen et al., 2017b) TPT 11 340.5 1.8 3 4 0.4
TPT-Bench (ours) TPT 48 397.2 65.3 9.6 2177 45.4
2.1 Target Person Tracking
@@
Target person tracking aims to identify and track a specific person, which is crucial for personalized HRI scenarios where the robot needs to engage with its user over extended periods (Islam et al., 2019; Singamaneni et al., 2024). Earlier TPT approaches often relied on visual tracking techniques, such as MOT or SOT, to track the target person. SOT-assisted methods initially identify the target person by selecting their bounding box in the first frame, then use an appearance model to search for the target within a nearby region (Chen et al., 2017b; Zhang et al., 2019). SOT techniques (Javed et al., 2022) aim to track an arbitrary object by simultaneously classifying and localizing it within a search region. However, due to the lack of prior knowledge about human appearance (e.g., from pre-trained models of people detection and person re-identification), these methods can easily track a wrong person in crowded environments, particularly after partial or total occlusions.
@@
Conversely, MOT techniques (Luo et al., 2021) assume prior knowledge of the tracked object, typically people. MOT techniques usually detect individuals using well-trained object detectors and track multiple people within the sensor’s field of view (FoV). Based on these well-tracked people, MOT-assisted TPT methods then identify the target person by selecting their identity  (Leigh et al., 2015; Repiso et al., 2020a). Such a solution is robust to short-term occlusions but still struggles with long-term occlusions since they often lack a target re-identification (ReID) capability after occlusions.
@@
To address these gaps, recent TPT methods typically combine MOT with a target ReID module. The MOT front-end serves two main purposes: 1) tracking people’s trajectories is essential for socially-aware path planning (Repiso et al., 2020a; Morales et al., 2017; Triebel et al., 2016), which helps the robot assist the target person while respecting proxemics with others, and 2) by re-identifying the target from tracked individuals, established person ReID techniques (Ye et al., 2021) can be utilized in the back-end module to recognize the target person after partial or total occlusions. With the advancement of deep learning, deep-learning-based (DL-based) ReID features have become dominant due to their robustness to viewpoint and illumination changes (Ye et al., 2021).
@@
To leverage the power of DL-based features for target ReID, several approaches have been proposed (Rollo et al., 2024; Koide et al., 2020; Ye et al., 2024). For example, Koide et al. (2020) introduced an online boosting classifier using low-level CNN features, while Rollo et al. (2024) represented the target feature as a distribution and updated it continuously with newly archived features using a damped exponential moving average. Moreover, Ye et al. (2024) observed a domain shift in pre-trained ReID features—those from video-surveillance data often fail to generalize to practical person-following environments. To address this, they proposed using online continual learning techniques to optimize ReID features by collecting valuable observations.
@@
This paper evaluates the most recent target-ReID methods (Rollo et al., 2024; Ye et al., 2024) and popular SOT trackers on TPT-Bench, as well as other related datasets like LaSOT (Fan et al., 2021) and Chen’s dataset (Chen et al., 2017b), which track a target person over extended periods. It highlights the more challenging long-term tracking scenarios presented by TPT-Bench.
2.2 Datasets related to Target Person Tracking
2.2.1 Datasets for Target Person Tracking
@@
Datasets for evaluating TPT are limited. The only public dataset (Chen et al., 2017a) is collected by teleoperating the robot following the target person, involving challenging scenarios such as quick multi-people crossings, illumination changes, and appearance variations. However, this dataset lacks challenging scenarios requiring person ReID, such as occlusion and similar appearances of distracting people. To mitigate this limitation, some methods propose custom-built datasets (Ye et al., 2024; Rollo et al., 2024; Koide et al., 2020), but their experimental scenarios are limited, involving only a few distracting people.
@@
Moreover, ROREAS (Gross et al., 2017) and ROREAS+ (Eisenbach et al., 2023) propose TPT datasets involving scenarios of stroke patient assistance in their daily rehabilitation. Although these include challenges such as frequent occlusions by multiple people and lighting changes, their scenarios are limited to stroke patient assistance and are private due to data protection laws.
2.2.2 Datasets for SOT and MOT
@@
Since TPT focuses on continuous online tracking of a target person, related datasets include MOT datasets (Milan, 2016; Manen et al., 2017; Dendorfer, 2020) and SOT datasets (Wu et al., 2013; Muller et al., 2018; Fan et al., 2021). SOT datasets, such as LaSOT (Fan et al., 2021) and TrackingNet (Muller et al., 2018), focus on tracking arbitrary objects, including people, across various scenarios like dance, surveillance, and sports. While these datasets present challenges like occlusion, cluttered backgrounds, and pose variations, they lack robot-egocentric sequences designed for personalized robot assistance. In such scenarios, the tracked person may frequently be occluded or disappear for long periods. Additionally, the robot’s perspective differs from that of surveillance or handheld cameras in that people are often partially observed with body parts, making it more challenging to track the person.
@@
MOT datasets, such as MOT20 (Dendorfer, 2020), are designed for multi-person tracking in crowded environments with frequent occlusions. However, they focus on surveillance footage, where occlusion patterns differ significantly from those seen in robot observations. These datasets also provide limited long-term tracking data for a target person. The JRDB dataset (Martin-Martin et al., 2021), the most relevant to our work, includes egocentric data with 3D LiDAR, IMU and visual data, collected by a teleoperated robot in crowded environments. However, it does not represent a long-term tracking scenario in the context of HRI. Moreover, it assumes people remain within the camera’s field of view (FoV) and does not address the need to re-identify individuals when they reappear after being occluded.
@@
A quantitative comparison of person-tracking datasets is presented in Table 1, with most metrics derived from a long-term tracking benchmark (Lukeźič et al., 2020). The TPT-Bench captures long-term robot companionship from an egocentric perspective in crowded, unstructured environments, resulting in more occlusions and longer disappearance periods compared to existing person-tracking datasets. Additionally, when a target reappears, its appearance may differ from its previous state in terms of pose, lighting, and scale, as shown in Figure 1, which presents additional challenges for target ReID. These difficulties are demonstrated in the experiments through a comprehensive comparison of current TPT algorithms on the TPT-Bench.
3 The TPT-Bench Dataset
@@
In this section, we first describe the data collection platform (Sec. 3.1.1) and introduce the scenarios (Sec. 3.1.2) collected in this dataset. We then describe the annotation process and standards (Sec. 3.2) and provide dataset statistics (Sec. 3.3) of the whole dataset, highlighting unique challenges present by the proposed dataset. In the end, we introduce the dataset organization (Sec. 3.4) and the development tools (Sec. 3.5).
3.1 Collecting Data
3.1.1 Platform